Please write the following with steps
Draw a graph of each function by first writing an equivalent exponential form and then making a table of values
I am only having trouble with writing an equivalent exponential form
y = log2(x plus 1)

Answer by stanbon(48569)

y = log2(x plus 1)
-----------------
The exponent is y
The base is 2
The inverse log or anti-log is x+1
--------------------------
x+1 = 2^y it the equivalent exponential equation.
================
Cheers,
Stan H.